OpenStudy (anonymous):

Prove the identity: (sinx)(cotx + cosx tanx) = cosx + sin^2x

OpenStudy (mertsj):

\[\sin x(\frac{\cos x}{\sin x}+\cos x \times \frac{\sin x}{\cos x})=\sin x(\frac{\cos x}{\sin x}+\sin x)=\cos x+\sin ^2x\]